# A Simple DropConnect Approach to Transfer-based Targeted Attack

### v1

@@ 4.3 Discussion
In this section, we compare MCD with related self-ensemble methods in more details.
While MCD is proposed from the perspective of mitigating perturbation co-adaptation, it turns out to be a self-ensemble method that constructs multiple new models at each iteration. Existing self-ensemble method [18] targets object detection task, which is different from ours. SE-ViT [28] is specifically designed for vision Transformer surrogate models; as shown in Sec. 5, its ASR is lower than MCD even when using ViT as surrogate model. More closely related are Ghost Network [22] and MUP [45], but are much outperformed by MCD (c.f. Sec. B.2 and Sec. 5.2).
We now analyze the effectiveness of MCD from model ensemble perspective. It has been hypothesized that an adversarial image that remains adversarial for multiple models is more likely to transfer to other models [26]. MCD generates independent random masks for each selected layer and also at each optimization iteration. In this sense, it brings in more randomness and further diversification than Ghost Network and MUP. Specifically, Ghost Network perturbs only skip connections for residual networks (like ResNet-50), while MUP masks unimportant parameters according to a predefined metric at each iteration. This observation is also in accordance with [18, 3, 25], which show that increasing the number of surrogate models generally enhances the transfer attack performance. On the other hand, from the perspective of ensemble techniques in machine learning, each variant model should also be informative about or useful to the targeted task of image classification [18, 3]. As empirically shown in Sec. 5.5, MCD achieves a good tradeoff in terms of model diversity and utility, thereby enhancing the overall attack performance.
In addition, MCD is simple to implement and is robust across a range of parameter choices. In particular, our implementation of MCD is to pick the same DropConnect probability, i.e., $p_{w}=p_{e}=p$, for selected layers. MCD can be applied to all layers in the model with little decrease of ASR, compared to our best practice that only selects the linear and normalization layers in this paper. See the experiments in Sec. 5 for more details.
@@ 5 Experiments
This section conducts extensive experiments to show the effectiveness of the proposed method. In particular, we consider both CNN- and Transformer-based classifiers.
@@ 5.1 Experimental Settings
We utilize the ImageNet-compatible dataset,https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset which served as the official dataset for the NIPS 2017 Attack Challenge. This dataset contains both ground-truth and targeted labels, making it well-suited for conducting experiments on targeted attacks.
@@
We adopt the $\ell_{\infty}$-norm as the constraint on perturbation, with budget $\epsilon=16/255$. The learning rate is chosen as $\alpha=2/255$. We use the Logit loss [47] as our objective function in Eq. 1. By default, we set the maximum number of optimization iterations to $1,000$ and the batch size to $32$ for all baseline methods, to ensure sufficient optimization. Furthermore, we notice that different attack methods may take different amounts of computations per iteration. We hence fix the same number of inferences for each optimization iteration to maintain fair comparisons across all attack methods.
@@
Surrogate and Target Models
Our experiments choose various models commonly used in the literature [2]. These include 1) CNN-based models: VGG-16 [32], ResNet-18 (RN18) [15], ResNet-50 (RN50) [15], DenseNet-121 (DN121) [17], Xception (Xcep) [4], MobileNet-v2 (MBv2) [31], EfficientNet-B0 (EFB0) [37], Inception ResNetv2 (IRv2) [36], Inception-v3 (Incv3) [35], and Inception-v4 (Incv4) [36]; and 2) Transformer-based models: ViT [9], LeViT [12], ConViT [10], Twins [5], and Pooling-based Vision Transformer (PiT) [16]. All the models are pre-trained on the ImageNet dataset [6] . Additionally, we include CLIP [29], trained on 400 million text-image pairs, to evaluate the transferability across different modalities. When using Transformer-based models, the input image is resized to 224$\times$224 to meet the model input requirement.
@@
We compare MCD with various existing transfer-based methods, including DI [44], RDI [48], SI [24], Admix [42], BSR [40], ODI [1], and CFM [2]. We also include two existing self-ensemble methods, namely, MUP (whose implementation only handles CNN layers) [45] and SE-ViT (which is specifically designed for vision Transformers) [28]. These methods are primarily used in combination with TI-FGSM [8] and MI-FGSM [7] during the optimization process. It is worth noting that some of these methods, namely, SI, BSR, Admix, CFM, MUP and SE-ViT, are implemented together with RDI, which have been reported to obtain higher transfer ASRs [2]. As previously mentioned, the attack methods are configured with an identical number of inferences per iteration, denoted as $S$. Specifically, we pick $S$ scaled copies in SI and in the inner loop of Admix, and $S$ transformed images for BSR . For other baselines, we perform $S$ forward-backward passes and use the average gradient on $x_{\mathrm{adv}}$ to update the adversarial example in each iteration. In the main experiment, we will set $S=5$; comparison of other choices is studied in Sec. 5.3.
MCD Setting
While the DropConnect probabilities can be chosen differently for the weight and bias parameters in a linear layer (or the transformation parameters in the normalization layer) and also across different layers, we simply pick the same probability for all selected parameters, that is, $p_{w}=p_{b}=p$, which greatly simplifies the implementation. Through our ablation study, we find that MCD performs well across a range of probabilities. For our experiments, we will select the following DropConnect probabilities: $0.05$ for ResNet-50, $0.02$ for Inception-v3, $0.04$ for DenseNet-121, $0.01$ for Vision Transformer, and $0.03$ for CLIP. By default, MCD applies DropConnect to all linear and normalization layers in the surrogate model.
Source: RN50 Source: DN121
DI 0.4 6.7 0.6 3.8 1.8 0.5 2.3 0.3 4.0 1.2 2.0 2.1 0.3 1.7
RDI 2.8 24.0 4.4 12.6 10.1 1.2 9.2 1.0 12.0 2.1 6.9 8.2 1.4 5.3
SI 8.0 42.9 7.7 25.1 23.3 3.7 18.4 3.9 21.5 4.4 10.2 14.3 2.1 9.4
BSR 6.8 42.4 7.7 25.3 21.9 2.3 17.7 2.7 21.6 2.7 10.6 11.9 1.2 8.5
Admix 7.2 43.4 5.9 22.7 19.4 4.2 17.1 4.2 33.8 4.0 18.0 19.1 3.9 13.8
ODI 15.5 49.3 11.8 31.6 35.1 5.6 24.8 8.5 36.3 8.4 19.8 26.7 4.9 17.4
MUP 7.0 48.3 8.2 30.7 26.5 3.6 20.7 5.2 38.9 5.0 23.2 20.1 4.1 16.1
CFM 17.3 65.8 14.6 47.5 39.9 7.9 32.2 10.4 49.0 8.3 30.1 30.4 5.5 22.3
MCD 34.5 76.7 27.1 60.1 57.1 15.7 45.2 28.0 68.5 23.2 53.8 53.6 14.1 40.2
Table 2: ASRs (%) against five transformer-based target models on the ImageNet-Compatible dataset. We also include the result on CLIP. All the attack methods are combined with MI-TI. The best results are shown in bold and the second best results are underlined.
@@ 5.2 Main Result
We first study the performance of MCD on the ImageNet-Compatible dataset. We employ ResNet-50, Inception-v3, DenseNet-121, and ViT as surrogate models, and evaluate the obtained adversarial examples on 16 target models.
Tab. 1 reports the ASRs of various attack methods on ten CNN-based target models. MCD achieves the best average ASR. Particularly, with Inception-v3 as surrogate model, the ASRs are increased by 14.8% and 20.9% for the challenging target models VGG16 and MBv2, respectively. When transferring from Transformer-based model ViT to CNN-based models, MCD again attains the best average ASR of 51.2%. We also report the results of self-ensemble methods MUP [45] and SE-ViT [28], and MCD clearly outperforms these two methods by a large margin.
In Tab. 2, we present experimental results of targeted attacks, where adversarial examples are generated using CNN-based models and are transferred to Transformer-based neural networks. This task is considered more challenging in the context of transfer-based attacks [27], as the ASRs in this case are still relatively low. Again, our method significantly improves the attack performance over existing methods: it increases the average ASR from 32.2% to 45.2% with ResNet-50 as surrogate model, and from 22.3% to 40.2% when the surrogate model is DenseNet-121.
Due to space limit, additional experimental results with different surrogate models can be found in Sec. B.4. We also visualize the heatmaps of some generated adversarial examples in Appendix A for qualitative comparison.
5.3 More Iterations and Inferences
We now study the performance given different total iterations and different numbers of inferences per iteration, which are denoted as $T$ and $S$, respectively. We use ResNet-50 as the surrogate model and evaluate how well the adversarial examples are transferred to the 16 target models.
Figure 4: Average ASRs with different iterations (denoted by $T$) and different numbers of inferences per iteration (denoted by $S$).
The results are reported in Fig. 3. Although existing methods may also benefit from additional optimization iterations, MCD and Admix have the best gains when $T$ increases, while MCD achieves a much higher ASR than Admix. With $S$ increasing, MCD can outperform CFM even at an early stage of the optimization process. We also depict Fig. 4 to ease the comparison of the gains of different methods when both $T$ and $S$ increase. As we observe, MCD benefits the most from an additional compute budget.
@@ 5.4 Ablation Study
In this section, we conduct an ablation study to investigate the impacts of 1) different types of layers where DropConnect is applied, and 2) different DropConnect probabilities.
@@
Table 3: Applying DropConnect to different types of layers. The reported ASRs(%) are averages over 16 different target models.
Different Types of Layers
We use ResNet-50 and ViT as surrogate models to analyze the impacts of layer types. For ResNet-50, we apply MCD to different combinations of Batch Normalization (BN) layer, Fully Connected (FC) layer, and Convolutional (Conv) layer. Similarly, we consider Layer Normalization (LN) and FC layers (including linear transformation layer in the attention layer) for ViT.
Tab. 3 presents the experimental results. Notably, simply applying MCD to all layers achieves equal or higher ASRs, compared with other baselines. The combination of BN (or LN) and FC layers performs the best, which validates our implementation in Sec. 4. For ResNet-50, applying DropConnect to Conv layers performs worse than BN layers. We conjecture that Conv layers generally have sparser weights and may be less affected by perturbation co-adaptation.
@@
We further investigate the impact of varying DropConnect probabilities, denoted as $p$, using ResNet-50 as a surrogate model. We run MCD with different probabilities ranging from $0.01$ to $0.09$ and report the average ASRs over 16 different models. Along different $p$, the mean of average ASRs is 66.3% with a standard deviation of 5.9%. The highest average ASR is 72.4%, achieved at $p=0.05$. Notably, with $p\in[0.03,0.07]$, MCD consistently outperforms other baselines by more than 2%, underscoring the stability of the proposed method across different choices. Due to space limit, the detailed results are given in Fig. 9 in the supplementary material.
5.5 Model Diversity and Utility
In this section, we show the diversity and utility of constructed models. Here diversity is measured by first computing the mean KL-divergence between the output distributions of each pair of model variants for an input and then averaging these values over the ImageNet-compatible dataset. Similarly, the utility is quantified by the mean top-1 accuracy over the model variants and the dataset. We use the same surrogate model ResNet-50 as in the ablation study.
The results are presented in Fig. 5. As DropConnect probability $p$ increases, the diversity among different variants improves while the utility decreases. As shown in Fig. 9 in the supplementary material, the average ASR using the same surrogate model increases with $p\in[0.01,0.05]$ and then decreases with $p\in[0.05,0.09]$, while the highest ASR is obtained at $p=0.05$. This verifies our discussion in Sec. 4.3 that the attack performance is affected by both diversity and utility. We conclude that the better performance of MCD is due to the improved diversification across the variants while keeping each variant sufficiently useful.
5.6 Attack Performance Against Defense Methods
Finally, we evaluate MCD against several defense mechanisms, including adversarial training methods adversarially trained ResNet-50 (advRN) [30] and Ensemble-Adversarial-Inception-ResNet-v2 (ensIR) [19], and also High-level representation Guided Denoiser (HGD) [23], Bit Depth Reduction (Bit) [14], JPEG compression [14], and R&P [43]. We utilize ResNet-50 as the surrogate model. For Bit, JPEG, and R&P, the target model is ResNet-18. As shown in Tab. 4, MCD  outperforms all other baselines. Notably, in the challenging scenario of ensIR and HGD defenses, our method surpasses the second best ASRs by 29.4% and 10.5%, respectively.
(a) Diversity
(b) Utility
Figure 5: ASRs(%) with increasing mask ratio.
@@ 5.6 Attack Performance Against Defense Methods
Method advRN ensIR JPEG Bit R&P HGD
DI 10.6 0.0 26.7 50.5 41.1 0.1
RDI 39.6 0.8 58.8 75.8 71.3 0.8
SI 61.8 9.4 75.1 80.7 78.8 2.1
Admix 68.9 5.7 76.9 81.4 78.6 2.4
BSR 60.4 3.0 76.8 85.2 83.4 2.6
ODI 58.6 5.1 71.7 73.8 76.1 0.3
CFM 84.1 13.8 88.2 91.5 89.3 15.2
MCD 88.2 43.2 91.2 92.2 92.7 25.7

### v2

@@ 3.3 Discussion
In this section, we compare RaPA with related self-ensemble methods in more details.
RaPA was proposed to reduce over-dependence on specific parameters, and it turns out to be a self-ensemble method that constructs multiple new models at each iteration. Existing self-ensemble method [19] targets object detection task, which is different from ours. SE-ViT [33] is specifically designed for vision Transformer surrogate models; as shown in Section 4, its ASR is lower than RaPA even when using ViT as surrogate model. More closely related are Ghost Network [25] and MUP [54], but are much outperformed by RaPA (c.f. Section D.1 and Section 4.2).
We now analyze the effectiveness of RaPA from model ensemble perspective. It has been hypothesized that an adversarial image that remains adversarial for multiple models is more likely to transfer to other models [30]. RaPA generates independent random masks for each selected layer and also at each optimization iteration. In this sense, it brings in more randomness and further diversification than Ghost Network, MUP and DWP. Specifically, Ghost Network perturbs only skip connections for residual networks (like ResNet-50), while MUP and DWP mask unimportant parameters according to a predefined metricor at each iteration. This observation is also in accordance with [19, 3, 29], which show that increasing the number of surrogate models generally enhances the transfer attack performance. On the other hand, from the perspective of ensemble techniques in machine learning, each variant model should also be informative about or useful to the targeted task of image classification [19, 3]. As empirically shown in Section D.4, RaPA achieves a good tradeoff in terms of model diversity and utility, thereby enhancing the attack performance.
@@ 4 Experiments
Source: RN50 Source: DN121
DI 0.4 6.7 0.6 3.8 1.8 0.5 2.3 0.3 4.0 1.2 2.0 2.1 0.3 1.7
RDI 2.8 24.0 4.4 12.6 10.1 1.2 9.2 1.0 12.0 2.1 6.9 8.2 1.4 5.3
SI 8.0 42.9 7.7 25.1 23.3 3.7 18.4 3.9 21.5 4.4 10.2 14.3 2.1 9.4
SIA 3.1 28.4 3.9 16.7 13.5 1.9 11.2 2.4 24.2 2.2 12.9 11.1 1.5 9.0
BSR 6.8 42.4 7.7 25.3 21.9 2.3 17.7 2.7 21.6 2.7 10.6 11.9 1.2 8.5
Admix 7.2 43.4 5.9 22.7 19.4 4.2 17.1 4.2 33.8 4.0 18.0 19.1 3.9 13.8
ODI 15.5 49.3 11.8 31.6 35.1 5.6 24.8 8.5 36.3 8.4 19.8 26.7 4.9 17.4
MUP 7.0 48.3 8.2 30.7 26.5 3.6 20.7 5.2 38.9 5.0 23.2 20.1 4.1 16.1
CFM 17.3 65.8 14.6 47.5 39.9 7.9 32.2 10.4 49.0 8.3 30.1 30.4 5.5 22.3
FTM 18.0 67.6 16.5 47.1 41.5 9.3 33.3 10.7 50.4 9.4 31.1 30.4 4.8 22.8
RaPA 33.8±1.0 75.4±1.8 27.6±0.2 59.5±0.4 57.3±0.7 15.6±0.1 45.0 27.8±0.1 69.4±0.6 23.5±0.2 53.1±0.6 54.0±0.1 14.1±0.0 40.3
Table 3: ASRs (%) against five Transformer-based target models on the ImageNet-Compatible dataset. All the attack methods are combined with MI-TI. The best results are shown in bold and the second best results are underlined.
This section empirically validates the effectiveness of our method, using both CNN- and Transformer-based models.
@@ 4.1 Experimental Settings
We utilize the ImageNet-compatible dataset [21], served as the official dataset for the NIPS 2017 Attack Challenge. This dataset contains both ground-truth and targeted labels, making it well-suited for targeted-attack.
@@
We adopt the $\ell_{\infty}$-norm as the constraint on perturbation, with budget $\epsilon=16/255$. The learning rate is chosen as $\alpha=2/255$. We use the Logit loss [56] as our objective function in Equation 1. By default, we set the maximum number of optimization iterations to $1,000$ and the batch size to $32$ for all baseline methods, to ensure sufficient optimization. Furthermore, we notice that different attack methods may take different amounts of computations per iteration. We hence fix the same number of inferences for each optimization iteration to maintain fair comparisons across all attack methods.
Surrogate and Target Models
Our experiments choose various models commonly used in the literature [2]. These include 1) CNN-based models: VGG-16 [38], ResNet-18 (RN18) [16], ResNet-50 (RN50) [16], DenseNet-121 (DN121) [18], Xception (Xcep) [4], MobileNet-v2 (MBv2) [37], EfficientNet-B0 (EFB0) [42], Inception ResNetv2 (IRv2) [39], Inception-v3 (Incv3) [40], and Inception-v4 (Incv4) [39]; and 2) Transformer-based models: ViT [10], LeViT [13], ConViT [6], Twins [5], and Pooling-based Vision Transformer (PiT) [17]. All the models are pre-trained on the ImageNet dataset [7] . Additionally, we include CLIP [35], trained on 400 million text-image pairs, to evaluate the transferability across different modalities. When using Transformer-based models, the input image is resized to $\times224$ to meet the model input requirement.
@@
We compare RaPA with various existing transfer-based methods, including DI [52], RDI [57], SI [28], Admix [48],SIA [49], BSR [46], ODI [1], and CFM [2]. We also include two existing self-ensemble methods, namely, MUP (whose implementation only handles CNN layers) [54] and SE-ViT (which is specifically designed for vision Transformers) [33]. These methods are primarily used in combination with TI-FGSM [9] and MI-FGSM [8] during the optimization process. It is worth noting that some of these methods, namely, SI, BSR, Admix, CFM, MUP and SE-ViT, are implemented together with RDI, which have been reported to obtain higher transfer ASRs [2]. As previously mentioned, the attack methods are configured with an identical number of inferences per iteration, denoted as $S$. Specifically, we pick $S$ scaled copies in SI and in the inner loop of Admix, and $S$ transformed images for BSR . For other baselines, we perform $S$ forward-backward passes and use the average gradient on $x_{\mathrm{adv}}$ to update the adversarial example in each iteration. In the main experiment, we will set $S=5$; comparison of other choices is studied in Section 4.3.
@@
RaPA Setting
While the DropConnect probabilities can be chosen differently for the weight and bias parameters in a linear layer (or the transformation parameters in the normalization layer) and also across different layers, we choose the same probability for all selected parameters, that is, $p_{w}=p_{b}=p$, which greatly simplifies the implementation. Through our ablation study, we find that RaPA performs well across a range of probabilities. For our experiments, we will select the following DropConnect probabilities: $0.05$ for ResNet-50, $0.02$ for Inception-v3, $0.04$ for DenseNet-121, $0.01$ for Vision Transformer, and $0.03$ for CLIP. By default, RaPA applies DropConnect to all linear and normalization layers in the surrogate model.
@@ 4.2 Main Result
We first study the performance of RaPA on the ImageNet-Compatible dataset. We employ ResNet-50, Inception-v3, DenseNet-121, and ViT as surrogate models, and evaluate the obtained adversarial examples on 16 target models.
Table 3 reports the experimental results when adversarial examples are generated using CNN-based models and transferred to Transformer-based neural networks. This task is considered more challenging in the context of transfer-based attacks [31], as the ASRs in this case are relatively low. Our method significantly improves the attack performance over existing methods: it increases the average ASR from 33.3% to 45.0% with ResNet-50 as surrogate model, and from 22.8% to 40.3% with DenseNet-121.
Table 4 reports the ASRs of various attack methods on ten CNN-based target models. RaPA achieves the best average ASR. Particularly, with Inception-v3 as surrogate model, the ASRs are increased by 14.6% and 20.7% for the challenging target models VGG16 and MBv2, respectively. When transferring from Transformer-based model ViT to CNN-based models, RaPA again attains the best average ASR 51.2%. We also report the results of self-ensemble methods MUP [54] and SE-ViT [33]. RaPA clearly outperforms these two methods by a large margin.
Additional experimental results can be found in Section D.3. We also visualize the heatmaps of some adversarial examples in Appendix C for qualitative comparison.
@@ 4.3 Ablation Study
In this section, we conduct an ablation study to investigate the impacts of 1) different types of layers where DropConnect is applied, 2) different DropConnect probabilities and 3) more iterations and inferences.
Different Types of Layers
We use ResNet-50 and ViT as surrogate models to analyze the impacts of layer types. For ResNet-50, we apply RaPA to different combinations of Batch Normalization (BN) layer, Fully Connected (FC) layer, and Convolutional (Conv) layer. Similarly, we consider Layer Normalization (LN) and FC layers (including linear transformation layer in the attention layer) for ViT.
Table 5 presents the experimental results. Notably, simply applying RaPA to all layers achieves equal or higher ASRs, compared with other baselines. The combination of BN (or LN) and FC layers performs the best, which validates our implementation in Section 3. For ResNet-50, applying DropConnect to Conv layers performs worse than BN layers. We conjecture that Conv layers have sparser weights and may be less affected by over-reliance issue. Applying DropConnect only to FC layers yields particularly low ASRs, as ResNet contains only a single FC layer.
@@
Table 5: Applying DropConnect to different types of layers. The reported ASRs(%) are averages over 16 target models.
@@
We investigate the impact of varying DropConnect probabilities $p$ using ResNet-50 as a surrogate model. RaPA is run with $p$ ranging from $0.01$ to $0.09$, and we report the average ASRs over 16 models in Sec. D.2. The mean ASR is 66.3% with a standard deviation of 5.9%, peaking at 72.4% when $p=0.05$. Notably, with $p\in[0.03,0.07]$, RaPA consistently outperforms baselines by over 2%, underscoring the stability of the proposed method across different choices.
More Iterations and Inferences
We study performance under different total iterations and numbers of inferences per iteration, denoted as $T$ and $S$, respectively. We use ResNet-50 as the surrogate model and evaluate how well adversarial examples transfer to the 16 target models.
Figure 3: Average ASRs with different iterations ($T$) and different numbers of inferences per iteration ($S$).
The results are reported in Figure 2. Although existing methods may also benefit from additional optimization iterations, RaPA and Admix have the best gains when $T$ increases, while RaPA achieves a much higher ASR than Admix. With $S$ increasing, RaPA can outperform CFM even at an early stage of the optimization process. We also depict Figure 3 to ease the comparison of the gains of different methods when both $T$ and $S$ increase. As we observe, RaPA benefits the most from an additional compute budget.
4.4 Attack Performance Against Defenses
We evaluate RaPA against several defenses: adversarially trained ResNet-50 (advRN) [36], Ensemble-Adversarial-Inception-ResNet-v2 (ensIR) [20], High-level representation Guided Denoiser (HGD) [27], Bit Depth Reduction (Bit) [15], JPEG compression [15], R&P [51], and Diffpure [34]. We utilize ResNet-50 as the surrogate model. For Bit, JPEG, and R&P, the target model is ResNet-18 and for Diffpure, the target model is ResNet-50. As shown in Table 6, RaPA outperforms all other baselines. Notably, against the strong defenses ensIR and HGD, RaPA exceeds the second-best ASRs by 29.4% and 10.5%, respectively.
@@ 4.4 Attack Performance Against Defenses
Method advRN ensIR JPEG Bit R&P HGD Diffpure
DI 10.6 0.0 26.7 50.5 41.1 0.1 0.0
RDI 39.6 0.8 58.8 75.8 71.3 0.8 0.0
SI 61.8 9.4 75.1 80.7 78.8 2.1 0.4
Admix 68.9 5.7 76.9 81.4 78.6 2.4 0.7
BSR 60.4 3.0 76.8 85.2 83.4 2.6 0.1
ODI 58.6 5.1 71.7 73.8 76.1 0.3 0.3
CFM 84.1 13.8 88.2 91.5 89.3 15.2 0.5
RaPA 88.2 43.2 91.2 92.2 92.7 25.7 4.0
@@
Attack ViT LeViT ConViT Twins PiT CLIP Avg.
RaPA 33.8 75.4 27.6 59.5 57.3 15.6 45.0
DSM 8.1 49.5 7.8 31.6 23.3 3.1 20.6
DSM-RaPA 50.8 84.2 42.2 74.6 72.6 25.1 58.3
SASD-WS 30.1 74.1 25.6 52.1 52.9 18.1 42.2
SASD-RaPA 42.9 79.5 35.3 63.0 63.1 25.9 51.6
Table 7: ASRs (%) of adversarial examples generated by ResNet50 against Transformer-based targets, comparing RaPA with training-based methods (DSM and SASD-WS) and their combinations. Note that when RaPA is combined with SASD-WS, we did not apply Weight Scaling.
4.5 Compatibility with Training-enhanced Frameworks
We further compare RaPA with two representative training-required approaches, DSM[53] and SASD-WS[50], both of which involve additional optimization to enhance surrogate models for better adversarial transferability. Specifically, DSM trains a surrogate model with dark knowledge extracted from a teacher model and enriched by mixing augmentation, which effectively boosts the model’s transferability through knowledge-driven soft supervision. Meanwhile, SASD-WS improves transferability via sharpness-aware self-distillation and weight scaling, refining the loss landscape and model generalization. As shown in Table 7, even under a fully training-free setting, RaPA already surpasses these training-dependent methods, exhibiting superior transferability and generalization. Furthermore, when integrated with such training-based schemes, RaPA continues to deliver consistent gains. For instance, combining with DSM increases the average ASR from 20.6% to 58.3%, highlighting its excellent compatibility and scalability with existing training-enhanced frameworks.